# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ Appendix A Technical lemmas
Consider a loss $\ell$ satisfying Assumption 3B. Then, we have
@@
Consider a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption 3A. Then, $\ell$ and $g$ are nonincreasing. Moreover, $\ell$ is strictly decreasing on $(-\infty,m)(\subseteq\mathbb{R})$ if $\ell$ has separation margin $m>0$; otherwise, $\ell$ is strictly decreasing on $\mathbb{R}$.
Proof.
@@
For the latter part, if nonincreasing $\ell$ has separation margin, $\ell(z)=0$ if and only if $z\geq m$. Then, we have $\ell\equiv 0$ on the interval $[m,\infty)\subseteq\mathbb{R}$ and $\ell>0$ on the interval $(-\infty,m)\subseteq\mathbb{R}$. On the latter interval, $\ell$ must be strictly decreasing because of its convexity. We can prove similarly for $\ell$ lacking separation margin. ∎
@@
Consider a loss $\ell$ satisfying Assumption 3B. Suppose that $\ell$ has separation margin $m>0$. Then, we have $\rho(\lambda)\leq m^{2}$ for any $\lambda\geq 1$.
Proof.
When $\ell$ has separation margin $m$ (see Definition 3), we have
@@
Suppose Assumption 2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumption 3C and let $\mathbf{u}\coloneqq\mathbf{u}_{1}+\mathbf{u}_{2}$ such that
@@
22 22 $\left\langle{\nabla L(\mathbf{w}_{k})},{\mathbf{u}_{2}}\right\rangle+\eta\|\nabla L(\mathbf{w}_{k})\|^{2}$
@@
$\leq\frac{2\|\mathbf{u}_{2}\|}{n}\sum_{i=1}^{n}\ell^{\prime}(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\left\langle{\mathbf{z}_{i}},{\mathbf{w}_{*}}\right\rangle+\eta C_{g}\cdot G(\mathbf{w}_{k})$ (Assumption 3C and $G(\mathbf{w}_{k})\geq 0$)
$\leq-2\gamma\|\mathbf{u}_{2}\|G(\mathbf{w}_{k})+\eta C_{g}\cdot G(\mathbf{w}_{k})$ (Assumption 2 and $G(\mathbf{w}_{k})\geq 0$)
@@
Telescoping the sum from 00 to $t-1$ and rearranging, we get
@@ Appendix B Proof of Theorem 5
Suppose Assumption 2 and consider (GD) with any stepsize $\eta>0$ under a Fenchel–Young loss $\ell$ that satisfies Assumption 3A. For $t\geq 1$, assume $G(\mathbf{w}_{k})\geq G_{\min}>0$ for $k\in[t]$, where $G(\mathbf{w})$ is defined in (7). Then, we have
@@
$=\left\langle{\mathbf{w}_{k}},{\mathbf{w}_{*}}\right\rangle+\frac{\eta}{n}\sum_{i=1}^{n}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\left\langle{\mathbf{w}_{*}},{\mathbf{z}_{i}}\right\rangle$ (use $g(\cdot)=-\ell^{\prime}(\cdot)$ by Lemma 9)
$\geq\left\langle{\mathbf{w}_{k}},{\mathbf{w}_{*}}\right\rangle+\gamma\eta G(\mathbf{w}_{k})$ (note $g(\cdot)\geq 0$ and Assumption 2)
@@
Consider a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption 3A. Then, for arbitrary $10<\bar{\varepsilon}<$ 1 and $\alpha$ defined in (3), we have
$g(\ell^{-1}(\varepsilon))\geq C_{\phi}\varepsilon^{\alpha}\quad\text{for $\varepsilon\in(0,\bar{\varepsilon})$},\quad\text{where}\quad C_{\phi}\coloneqq\frac{g(\ell^{-1}(\bar{\varepsilon}))}{\bar{\varepsilon}^{\alpha}}.$
Proof.
Choose any $\varepsilon_{0}>0$. For $\varepsilon\in(0,\varepsilon_{0})$, we can invert to get $z\equiv\ell^{-1}(\varepsilon)$ because $\ell$ is strictly decreasing on $\ell^{-1}((0,1))$ (by Lemma 9) and hence invertible. Note that
@@
for $z\in\mathcal{I}$, where we used the first-order optimality $-z=\phi^{\prime}(x)$ of the convex conjugate at the last identity. Since $\phi$ is twice continuously differentiable and $\phi^{\prime\prime}>0$ on the interval $(0,1)$ by Assumption 3A, we can apply the inverse function theorem to have
@@
for $z\in\mathcal{I}$. Since $\phi$ is twice continuously differentiable, we can apply the inverse function theorem once again to get ′′$\ell^{\prime\prime}(z)$ for $z\in\mathcal{I}$, and hence $g$ is differentiable on $\mathcal{I}$.
In addition, $\ell$ is continuously differentiable with non-degenerate derivative at $z\in\mathcal{I}$ because $\ell$ is strictly decreasing on $\mathcal{I}$. From these observations, we can see that $g(\ell^{-1}(\cdot))\eqqcolon f(\cdot)$ is nondecreasing on $\mathcal{I}$ and differentiable on $\mathcal{I}$ (because it is the composition of two nondecreasing and differentiable functions $g$ and $\ell^{-1}$). Now we can apply Lemma 13 to this $f$. Let us compute the exponent $\alpha_{\varepsilon}$ defined in (9). By the differentiability and non-degenerate derivative of $\ell$ on $\mathcal{I}$, we can apply the inverse function theorem on $\ell$ to have
@@
=(A)${=}\frac{[\mu\phi^{\prime}(\mu)-\phi(\mu)]\cdot\frac{1}{\phi^{\prime\prime}(\mu)}}{\mu^{2}}$ ($\mu\equiv(\phi^{*})^{\prime}(\bar{z})$)
@@
$[\phi^{\prime\prime}(\mu)]\cdot[(\phi^{*})^{\prime\prime}(\bar{z})]=1$
@@
by the definition of the convex conjugate and Crouzeix’s identity. Now we are ready to apply Lemma 13, which yields
$g(\ell^{-1}(\varepsilon))\geq C_{\phi}\varepsilon^{\alpha}\quad\text{for $\varepsilon\in(0,\bar{\varepsilon})$},$
@@
Proof of Theorem 5..
For $k\in[T-1]$, if $L(\mathbf{w}_{k})>\varepsilon$, there exists $i\in[n]$ such that $\ell(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)>\varepsilon/n$. Then, we have for this $i\in[n]$,
$\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle\leq\ell^{-1}(\varepsilon/n)$ ($\ell$ is strictly decreasing when $\ell>0$ by Lemma 9)
$\implies$ $g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\geq g(\ell^{-1}(\varepsilon/n)).$ ($g$ is nonincreasing by Lemma 9)
@@
$G(\mathbf{w}_{k})=\frac{1}{n}\sum_{j\in[n]}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{j}}\right\rangle)\geq\frac{1}{n}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\geq\frac{1}{n}g\left(\ell^{-1}\left(\frac{\varepsilon}{n}\right)\right)$
holds while $L$ is $\varepsilon$-suboptimal, that is, $L(\mathbf{w}_{k})>\varepsilon$.
Next, fix $\mathbf{w}_{0}=\mathbf{0}$. By Lemma 9, we can use Lemma 15. By Lemmas 12 and 15, we can take $G_{\min}=g(\ell^{-1}(\varepsilon/n))/n$ and have
$\frac{\gamma\eta g\left(\ell^{-1}\left(\frac{\varepsilon}{n}\right)\right)T}{n}$ $\leq\left\langle{\mathbf{w}_{T}},{\mathbf{w}_{*}}\right\rangle-\left\langle{\mathbf{w}_{0}},{\mathbf{w}_{*}}\right\rangle$ (Lemma 12)
@@
$\leq\|\mathbf{w}_{T}\|$ (the Cauchy-Schwarz inequality with $\|\mathbf{w}_{*}\|=1$)
$\leq\frac{4\sqrt{\rho(\gamma^{2}\eta T)}+\eta C_{g}}{\gamma}.$ (Lemma 15)
By reorganizing and applying Lemma 14, we leverage the primal-dual relationship to have
$T\leq\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\cdot\frac{1}{g\left(\ell^{-1}\left(\frac{\varepsilon}{n}\right)\right)}\leq\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\cdot\frac{1}{C_{\phi}\left(\frac{\varepsilon}{n}\right)^{\alpha}},$
where $C_{\phi}$ is defined in Lemma 14. Therefore, $L$ is $\varepsilon$-suboptimal after at most
$\frac{n^{1+\alpha}}{C_{\phi}\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta t)}}{\eta}+C_{g}\right)\varepsilon^{-\alpha}$
iterations. Finally, we verify that $C_{\phi}$ defined in Lemma 14 matches (3). By introducing $\bar{z}$ as the dual of $\bar{\mu}$ such that
@@
$\bar{\varepsilon}$ $=\ell(g^{-1}(\bar{\mu}))$ ($g$ is invertible when $10<g(\cdot)<$ 1)

### v2

@@ Appendix A Technical lemmas
Consider a loss $\ell$ satisfying Assumption˜3B. Then, we have
@@
Consider a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption˜3A. Then, $\ell$ and $g$ are nonincreasing. Moreover, $\ell$ is strictly decreasing on $(-\infty,m)(\subseteq\mathbb{R})$ if $\ell$ has separation margin $m>0$; otherwise, $\ell$ is strictly decreasing on $\mathbb{R}$.
Proof..
@@
For the latter part, if nonincreasing $\ell$ has separation margin, $\ell(z)=0$ if and only if $z\geq m$. Then, we have $\ell\equiv 0$ on the interval $m,\infty)\subseteq\mathbb{R}$ and $\ell>0$ on the interval $(-\infty,m)\subseteq\mathbb{R}$. On the latter interval, $\ell$ must be strictly decreasing because of its convexity. We can prove similarly for $\ell$ lacking separation margin. ∎
@@
Consider a loss $\ell$ satisfying Assumption˜3B. Suppose that $\ell$ has separation margin $m>0$. Then, we have $\rho(\lambda)\leq m^{2}$ for any $\lambda\geq 1$.
Proof..
When $\ell$ has separation margin $m$ (see Definition˜3), we have
@@
Suppose Assumption˜2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumption˜3C and let $\mathbf{u}\coloneqq\mathbf{u}_{1}+\mathbf{u}_{2}$ such that
@@
$2\left\langle{\nabla L(\mathbf{w}_{k})},{\mathbf{u}_{2}}\right\rangle+\eta\|\nabla L(\mathbf{w}_{k})\|^{2}$
@@
$\leq\frac{2\|\mathbf{u}_{2}\|}{n}\sum_{i=1}^{n}\ell^{\prime}(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\left\langle{\mathbf{z}_{i}},{\mathbf{w}_{*}}\right\rangle+\eta C_{g}\cdot G(\mathbf{w}_{k})$ (Assumption˜3C and $G(\mathbf{w}_{k})\geq 0$)
$\leq-2\gamma\|\mathbf{u}_{2}\|G(\mathbf{w}_{k})+\eta C_{g}\cdot G(\mathbf{w}_{k})$ (Assumption˜2 and $G(\mathbf{w}_{k})\geq 0$)
@@
Telescoping the sum from 0 to t1t-1 and rearranging, we get
@@ Appendix B Proof of Theorem 5
Suppose Assumption˜2 and consider (GD) with any stepsize $\eta>0$ under a Fenchel–Young loss $\ell$ that satisfies Assumption˜3A. For $t\geq 1$, assume $G(\mathbf{w}_{k})\geq G_{\min}>0$ for all $k\in[0,t-1]$, where $G(\mathbf{w})$ is defined in (6). Then, we have
@@
$=\left\langle{\mathbf{w}_{k}},{\mathbf{w}_{*}}\right\rangle+\frac{\eta}{n}\sum_{i=1}^{n}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\left\langle{\mathbf{w}_{*}},{\mathbf{z}_{i}}\right\rangle$ (use $g(\cdot)=-\ell^{\prime}(\cdot)$ by Lemma˜8)
$\geq\left\langle{\mathbf{w}_{k}},{\mathbf{w}_{*}}\right\rangle+\gamma\eta G(\mathbf{w}_{k})$ (note $g(\cdot)\geq 0$ and Assumption˜2)
@@
Consider a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption˜3A. Then, for arbitrary $0<\bar{\varepsilon}<1$ and $\alpha$ defined in (3), we have
[…]
Proof..
Choose any $\varepsilon_{0}>0$. For $\varepsilon\in(0,\varepsilon_{0})$, we can invert to get $z\equiv\ell^{-1}(\varepsilon)$ because $\ell$ is strictly decreasing on $\ell^{-1}((0,1))$ (by Lemma˜8) and hence invertible. Note that
@@
for $z\in\mathcal{I}$, where we used the first-order optimality $z=\phi^{\prime}(x)$ of the convex conjugate at the last identity. Since $\phi$ is twice continuously differentiable and $\phi^{\prime\prime}>0$ on the interval $(0,1)$ by Assumption˜3A, we can apply the inverse function theorem to have
@@
for $z\in\mathcal{I}$. Since $\phi$ is twice continuously differentiable, we can apply the inverse function theorem once again to get $\ell^{\prime\prime}(z)$ for $z\in\mathcal{I}$, and hence $g$ is differentiable on $\mathcal{I}$.
In addition, $\ell$ is continuously differentiable with non-degenerate derivative at $z\in\mathcal{I}$ because $\ell$ is strictly decreasing on $\mathcal{I}$. From these observations, we can see that $g(\ell^{-1}(\cdot))\eqqcolon f(\cdot)$ is nondecreasing on $\mathcal{I}$ and differentiable on $\mathcal{I}$ (because it is the composition of two nondecreasing and differentiable functions $g$ and $\ell^{-1}$). Now we can apply Lemma˜12 to this $f$. Let us compute the exponent $\alpha_{\varepsilon}$ defined in (8). By the differentiability and non-degenerate derivative of $\ell$ on $\mathcal{I}$, we can apply the inverse function theorem on $\ell$ to have
@@
$\overset{\text{(A)}}{=}\frac{[\mu\phi^{\prime}(\mu)-\phi(\mu)]\cdot\frac{1}{\phi^{\prime\prime}(\mu)}}{\mu^{2}}$ ($\mu\equiv(\phi^{*})^{\prime}(\bar{z})$)
@@
$\phi^{\prime\prime}(\mu)]\cdot[(\phi^{*})^{\prime\prime}(\bar{z})]=1$
@@
by the definition of the convex conjugate and Crouzeix’s identity. Now we are ready to apply Lemma˜12, which yields
[…]
@@
Proof of Theorem˜5..
For a fixed $k\in[T-1]$, if we have $L(\mathbf{w}_{k})>\varepsilon$, there exists $i\in[n]$ such that $\ell(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)>\varepsilon$. Then, we have for this specific $i\in[n]$,
$\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle<\ell^{-1}(\varepsilon)$ ($\ell$ is strictly decreasing when $\ell>0$ by Lemma˜8)
$\implies$ $g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\geq g(\ell^{-1}(\varepsilon)).$ ($g$ is nonincreasing by Lemma˜8)
@@
$G(\mathbf{w}_{k})=\frac{1}{n}\sum_{j\in[n]}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{j}}\right\rangle)\geq\frac{1}{n}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i}}\right\rangle)\geq\frac{1}{n}g\left(\ell^{-1}\left(\varepsilon\right)\right)$ (9)
holds while $L(\mathbf{w}_{k})$ is $\varepsilon$-suboptimal, that is, $L(\mathbf{w}_{k})>\varepsilon$.
Next, fix $\mathbf{w}_{0}=\mathbf{0}$ and consider the case where $L(\mathbf{w}_{k})>\varepsilon$ holds for all $k\in[T-1]$. By Lemma˜8, we can use Lemma˜19. By Lemmas˜11 and 19, we can take $G_{\min}=g(\ell^{-1}(\varepsilon))/n$ (noting (9)) and have
$\frac{\gamma\eta g\left(\ell^{-1}\left(\varepsilon\right)\right)T}{n}$ $\leq\left\langle{\mathbf{w}_{T}},{\mathbf{w}_{*}}\right\rangle-\left\langle{\mathbf{w}_{0}},{\mathbf{w}_{*}}\right\rangle$ (Lemma˜11)
@@
$\leq\|\mathbf{w}_{T}\|$ (the Cauchy–Schwarz inequality with $\|\mathbf{w}_{*}\|=1$)
$\leq\frac{4\sqrt{\rho(\gamma^{2}\eta T)}+\eta C_{g}}{\gamma}.$ (Lemma˜19)
By reorganizing and applying Lemma˜13, we leverage the primal-dual relationship to have
$T\leq\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\cdot\frac{1}{g\left(\ell^{-1}\left(\varepsilon\right)\right)}\leq\frac{n}{\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\cdot\frac{1}{C_{\phi}\varepsilon^{\alpha}},$
where $C_{\phi}$ is defined in Lemma˜13. Therefore, $L(\mathbf{w}_{k})$ is $\varepsilon$-suboptimal after at most
$\frac{n}{C_{\phi}\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta t)}}{\eta}+C_{g}\right)\varepsilon^{-\alpha}\qquad(\eqqcolon T(\varepsilon))$
iterations. That is, if $T>T(\varepsilon)$, the gradient lower bound (9) must be violated at some $t\in[T]$, and for this $t$, we achieve $L(\mathbf{w}_{t})\leq\varepsilon$.
Finally, we verify that $C_{\phi}$ defined in Lemma˜13 matches (3). By introducing $\bar{z}$ as the dual of $\bar{\mu}$ such that
@@
$\bar{\varepsilon}$ $=\ell(g^{-1}(\bar{\mu}))$ ($g$ is invertible when $0<g(\cdot)<1$)